OpenStudy (anonymous):

Reduce: (32 x^7 y^5 z^9 a^2)^(1/3) I got: (4z^3 x^2 y)(a^2 xy)^(1/3) as my answer. Am I wrong?

OpenStudy (anonymous):

\[\sqrt[3]{{32}x ^{7}y ^{5}z ^{9}a ^{2}}\]

OpenStudy (anonymous):

Thats a cubed root btw.

OpenStudy (anonymous):

\[4z ^{3}x ^{2}y \sqrt[3]{a ^{2}xy}\] is what i got for my answer

jhonyy9 (jhonyy9):

4 not can be because 4*4*4=64

jhonyy9 (jhonyy9):

cuberoot32 = 2cuberoot4

jhonyy9 (jhonyy9):

x7 =xcuberootx4

OpenStudy (anonymous):

Ah yeah I see what I did wrong

jhonyy9 (jhonyy9):

y5=ycuberooty2

jhonyy9 (jhonyy9):

z9=z2cuberootz

jhonyy9 (jhonyy9):

a2 reamind inside in this form

OpenStudy (anonymous):

but the cuberoot of z^9 is z^3 isnt it?

OpenStudy (anonymous):

z^3 x z^3 x z^3 = z^9

jhonyy9 (jhonyy9):

not is because this is cuberoot not squareroot

jhonyy9 (jhonyy9):

no what you have wr

jhonyy9 (jhonyy9):

e is x27

jhonyy9 (jhonyy9):

sorry z27

jhonyy9 (jhonyy9):

right ?

OpenStudy (anonymous):

but when you multuply you add the exponents

OpenStudy (anonymous):

you dont multiply the exponents

jhonyy9 (jhonyy9):

when you have exponent of exponent you need multiply exponentes

jhonyy9 (jhonyy9):

how cuberoot8 =2

jhonyy9 (jhonyy9):

is 2*2*2

jhonyy9 (jhonyy9):

so

OpenStudy (anonymous):

2 x 2 x 2 = 8 z x z x z x z x z x z x z x z x z = z^9

OpenStudy (anonymous):

2^3 = 8 (z^3)^3 = z^9

jhonyy9 (jhonyy9):

yes but z9 =(z3)3 = z3*z3*z3 =z27

OpenStudy (anonymous):

No you dont multiply exponents, you add them

OpenStudy (anonymous):

you just said z^9 = z^27 in your statement lol

jhonyy9 (jhonyy9):

so but there is cuberootz9

jhonyy9 (jhonyy9):

z9=z3*z3*z3 cuberootz3=z so cuberootz9=z*z*z=z3

OpenStudy (anonymous):

Yeah that's what I said

jhonyy9 (jhonyy9):

ok now is right ALL ?

OpenStudy (anonymous):

Yeah, just remember when you multiply, you add the exponents, and when you divide, you subtract the exponents.

jhonyy9 (jhonyy9):

can solve now your exercise ?

OpenStudy (anonymous):

Yeah I did, the problem with my original answer was the 4 on the outside. Should have been a 2 on the outside and left a 4 inside the squareroot

jhonyy9 (jhonyy9):

cuberoot

OpenStudy (anonymous):

Yeah thats what I meant my bad

jhonyy9 (jhonyy9):

ok bye

OpenStudy (anonymous):

Thanks :)

jhonyy9 (jhonyy9):

my pleasure

jhonyy9 (jhonyy9):

good luck